### v2

@@ Abstract
Autonomous agents that execute human tasks by controlling computers can enhance human productivity and application accessibility. However, progress in this field will be driven by realistic and reproducible benchmarks. We present AndroidWorld, a fully functional Android environment that provides reward signals for 116 programmatic tasks across 20 real-world Android apps. Unlike existing interactive environments, which provide a static test set, AndroidWorld dynamically constructs tasks that are parameterized and expressed in natural language in unlimited ways, thus enabling testing on a much larger and more realistic suite of tasks. Reward signals are derived from the computer’s system state, making them durable across task variations and extensible across different apps.
To demonstrate AndroidWorld’s benefits and mode of operation, we introduce a new computer control agent, M3A. M3A can complete 30.6% of AndroidWorld’s tasks, leaving ample room for future work. Furthermore, we adapt a popular desktop web agent to work on Android, which we find to be less effective on mobile, suggesting future research is needed to achieve universal, cross-domain agents. Finally, we conduct a robustness analysis by testing M3A against a range of task variations on a representative subset of tasks, demonstrating that variations in task parameters can significantly alter a task’s complexity and, consequently, an agent’s performance, highlighting the importance of testing agents under diverse conditions. AndroidWorld and the experiments in this paper are available at https://github.com/google-research/android_world.
@@ 1 Introduction
Autonomous agents that interpret human natural language instructions and operate computing devices can provide enormous value by automating repetitive tasks, augmenting human intelligence, and accomplishing complex workflows. The enthusiasm for building such agents is evident by the growing number of computer control agent systems [45, 10, 78, 27, 25, 19] and code repositories [16, 1, 2, 63, 66]. Yet, most existing approaches evaluate agents with proxy metrics, by comparing an agent’s trajectory to a previously collected human demonstration [32, 5, 10, 45, 70, 76, 37, 74, 68]. This kind of measurement is not representative of real-world performance [45, 78] because it does not reflect that there are usually multiple correct paths to solve a task, environments may behave non-deterministically (e.g., the environment freezes), and agents can dynamically learn from mistakes to correct their actions [50, 35, 31, 43]. Towards more accurate evaluation, emerging testing environments [71, 81, 27, 11, 29, 65] place agents in interactive environments and measure their task completion rates across a suite of tasks.
Reward signals are quantitative metrics that indicate functional correctness of a task, i.e. is the stated goal achieved? For example, for the task “Send a text message to Jane confirming I’ll be there”, a positive reward indicates the relevant message has been sent. Yet, real-world environments (apps and websites), unlike simulations [54, 51] and games [42, 52, 60, 62, 53], do not naturally provide explicit reward signals. The internals of most apps and websites are hidden from developers making it difficult to program rewards. Additionally, such systems are seemingly infinitely expansive, spanning millions of continuously-changing apps and websites. All these factors make it hard to establish reliable and durable reward functions.
One way to obtain reliable reward signals is to leverage human judgment [45, 78, 43, 26] to conduct one-off evaluations on a diverse suite of tasks; however, each evaluation run requires both significant time and costly manual review. LLM-based “judges” used to evaluate natural language generation models  [8, 79, 36] are emerging as a solution for evaluating agents in dynamic real-world environments [12, 39], and while showing promise also for computer control agents [43, 19] they remain imperfect evaluators. As a result, existing testing environments for computer control agents that provide automated reward signals are limited in their real-world diversity and scale. They cover mostly single-app tasks on a small set of domains (1–6 websites [71, 81, 27]) and evaluate agents on static test sets [65, 41], instead of across varied conditions and inputs which are found in real-world scenarios.
@@
To tackle these challenges, we introduce and release AndroidWorld, a robust agent environment for Android that offers realism and scalability for developing and evaluating computer control agents. AndroidWorld is an open environment consisting of a fully functional Android OS, allowing access to millions of Android apps and the Web. Crucially, it provides a highly reproducible task suite for developing and evaluating computer control agents, consisting of 116 tasks across 20 apps. Similarly to the popular MiniWoB++ [49] benchmark, each task is dynamically instantiated using randomly-generated parameters, challenging agents with millions of unique task goals and conditions. While MiniWob++ consists of simple, synthetic websites, AndroidWorld leverages real-world Android applications. To ensure that reward signals are durable when using real applications, AndroidWorld’s key insight is to leverage the extensive and consistent state management capabilities of the Android operating system, using the same mechanisms that the apps themselves utilize. The dynamic nature of AndroidWorld’s tasks is not only a test of adaptability for evaluation purposes, but it can also spawn new research on online learning algorithms for computer control agents.
AndroidWorld is a lightweight environment, requiring only 2 GB of memory and 8 GB of disk space, and is designed with convenience in mind. It connects agents to Android OS by leveraging the Python library AndroidEnv [58] to connect to the freely available Android Emulator. [^1] The Android Emulator is packaged as part of Android Studio, which can be downloaded from https://developer.android.com/studio AndroidWorld is highly extensible, making it easy to add new tasks and even new benchmarks, which we demonstrate by integrating the MiniWoB++ [49, 33] benchmark.
To demonstrate AndroidWorld’s usefulness as a benchmark, we build and release a new multi-modal agent, M3A (Multimodal Autonomous Agent for Android), and establish state-of-the-art results on AndroidWorld. We analyze M3A’s performance using both multimodal and text-only input, and we observe that while multimodal perception can improve performance in some cases, it generally does not outperform the text-only approach. On AndroidWorld, M3A achieves a 30.6% success rate, which surpasses that of a web agent adapted for Android but remains significantly lower than the human success rate of 80.0%. In pursuit of building robust computer control agents, our study includes comprehensive tests under varied real-world conditions, demonstrating significant performance variations primarily driven by changes in intent parameters.
We make the following contributions: (i) the creation of a new, highly diverse and realistic computer control agent environment; (ii) establishment of benchmark performance with a state-of-the-art multimodal agent, and (iii) a careful analysis demonstrating the need to evaluate agents across multiple test sets and conditions due to the inherent stochasticity in both models and environments.
@@ 2 Related Work
Table 1 compares existing evaluation environments for autonomous agents.
@@ 2.1 Interactive evaluation environments
Effective evaluation of autonomous agents requires environments that not only mimic real-world scenarios but also provide immediate reward signals upon successful task completion [45, 10, 3, 46, 6]. MiniWoB++ [49, 35] offers a lightweight framework of small, synthetic HTML pages with parameterized tasks which allow for unlimited task variability. WebShop [71] provides a simulated e-commerce environment, whereas WebArena [81] and VisualWebArena [27] consist of simulated websites across up to four domains. WorkArena [11] consists of 29 tasks for enterprise software. OSWorld [65] provides an interface and programmatic rewards across nine apps for desktop OSes. GAIA [41] is a static dataset that tests an agent’s ability to interact with live web environments. MMInA [77] is a multihop and multimodal benchmark designed to evaluate agents for compositional Internet tasks. B-MoCA [29] is the only interactive testing environment that exists today for mobile. It supports 31 tasks whose success is determined by comparing an agent’s execution with logs from human demonstrations using simple regex expressions. Unlike MiniWoB++, these frameworks use a fixed set of evaluation tasks with static initial conditions and goals, limiting adaptability. As has been observed in other domains such as reinforcement learning [20, 44, 9], the current output of foundation models is stochastic and highly sensitive to their inputs [23, 38, 34, 47], motivating the need to test them under varying conditions.
Reward signal construction is crucial when comparing interactive benchmarking environments. MiniWoB++  [49] dynamically constructs tasks, embedding rewards directly in custom HTML and JS code, a method specific to web-based tasks and less adaptable to other environments. WebArena [81], VisualWebArena [27] and MMInA [77] check trajectories for specific URLs and examine result pages for expected strings or images. Information retrieval tasks compare returned answers against a ground truth. WebShop [71] identifies specific products by matching their attributes against predefined gold attributes. B-MoCA [29] uses regular expressions to match log outputs. These types of reward are application specific and time dependent. OSWorld [65] scales better by inspecting device states or querying the cloud to compute rewards, thus offering robust functions adaptable to content and app updates, and shareable across multiple tasks. AndroidWorld enhances OSWorld’s approach by dynamically constructing the start states and varying the task goals in unlimited ways.
@@
Other studies leverage human evaluation [45, 78, 4] for tasks where automatic evaluation is not available. Lastly, emerging research [43, 19] explores the potential of multimodal models to generalize agent evaluations to new settings, though this area requires further research to achieve accuracy comparable to manually coded rewards.
2.2 Static datasets for automation
@@
Datasets derived from human interactions provide proxy metrics that correlate with real-world agent performance [32, 5, 10, 45]. On mobile platforms, AitW [45], PixelHelp [32], UGIF [59], and MoTIF [5] consist of demonstrations across Android apps and mobile websites, with screens often represented via accessibility trees. In contrast, desktop web environments typically utilize the DOM for representing website content, with Mind2Web [10], OmniAct [24] and others, across various desktop websites. Mobile-based datasets frequently involve more complex actions, such as scrolling, which are not as useful in DOM-based desktop interactions where the entire action space is readily accessible. Additionally, API-centric datasets like API-Bank [30], ToolTalk [13], and ToolBench [67] assess agents’ capabilities to manipulate computer systems via APIs.
@@ 2.3 Interactive agents
Prior to today’s foundation models, traditional approaches to developing user interface-operating agents primarily used reinforcement learning and behavioral cloning to simulate interactions like mouse clicks and keyboard typing [32, 35, 17, 22]. More recent work tends to leverage off-the-shelf foundational models [55, 56, 57] with in-context learning (ICL) and fine-tuning applied to mobile [45, 21, 61, 68, 76, 4, 75], desktop web [78, 10, 81, 27, 7, 28], and desktop OS [64, 73, 65]. Recent work explores agents that reflect on system state [50, 72, 40] by leveraging exploration, self-evaluation, and retry-capabilities enabling continual learning and adaptation [31, 70, 43, 64, 15].
@@
B-MoCA  12    31    1.9 regex Android (Apps+Web)
@@ 3.2 The observation and action space
AndroidWorld provides an interface for agents to receive observations and execute actions on Android. It uses AndroidEnv [58] and the Android Device Bridge to facilitate interaction between Android and the agent. The observation space consists of a full-resolution screenshot and a UI tree representation developed for accessibility purposes. The action space is similar to that which humans use, consisting of gestures (i.e., tapping, long-press, and swiping), typing, and navigation buttons (i.e., go home and go back). In addition to these naturalistic actions, AndroidWorld exposes a limited set of function calling APIs, such as send_text_message, to help agents accomplish goals. App. B provides more details on the observation format and action space.
@@ 3.3 Reproducible and parameterized tasks
Figure 2: Annotators performed the tasks assigned to them, assigned a difficulty level (2(a)) and selected relevant category tags from a predefined list (2(b)). They also estimated the number of steps required to complete each task (2(b)), using the action space available to an agent.
AndroidWorld consists of a suite of 116 tasks, spread across 20 diverse applications (see App. C for more details). These tasks simulate practical, everyday activities, including note-taking, scheduling appointments, communicating through messaging, and interacting with system utilities. The suite consists of open-source apps and built-in Android system apps, such as Settings and Contacts. As rated by humans, the tasks vary in difficulty, duration, and categories (Figure 2).
To achieve a high degree of reproducibility in real-world scenarios, AndroidWorld precisely controls the OS and app states in several ways. The Android OS is fixed, consisting of a Pixel 6 emulator running Android 13 with a fixed time on the date October 15th, 2023. All applications in AndroidWorld are fully-functional, open-source apps, besides the OS-level apps included with Android. For the open-source apps, AndroidWorld maintains a constant environment by installing a fixed version of each app, acquired from F-Droid.  [^4] https://f-droid.org/. OS-level apps’ versions are determined by the Android OS, which is also fixed. AndroidWorld only utilizes apps that do not require login/authentication, can function offline, and store their application data on device.
In addition to managing the states of apps and operating systems, AndroidWorld precisely defines and controls the state during task execution. Each task has its own unique setup, reward determination logic, and teardown procedures (see App. C.2 and C.3 for more details), ensuring a fully reproducible suite of tasks. This careful state management not only ensures reproducibility but also allows for easy customization of tasks through parametrization. Task parameters, initialized randomly at the start of each task based on a controlled random seed, dictate the initial state and influence reward outcomes. Similarly to MiniWoB++ [49, 33], AndroidWorld consists of a practically infinite set of varying initial conditions and success criteria. This approach provides more granular analyses of agents’ adaptability — a vital attribute for real-world deployment. Beyond testing agent robustness, the dynamic construction of tasks supports the use of online learning methodologies, particularly reinforcement learning [49, 33, 22, 17]. It also simplifies the generation of distinct train/test datasets, facilitating supervised learning experiments [22, 48, 14].

### v3

@@ Abstract
Autonomous agents that execute human tasks by controlling computers can enhance human productivity and application accessibility. However, progress in this field will be driven by realistic and reproducible benchmarks. We present AndroidWorld, a fully functional Android environment that provides reward signals for 116 programmatic tasks across 20 real-world Android apps. Unlike existing interactive environments, which provide a static test set, AndroidWorld dynamically constructs tasks that are parameterized and expressed in natural language in unlimited ways, thus enabling testing on a much larger and more realistic suite of tasks. To ensure reproducibility, each task includes dedicated initialization, success-checking, and tear-down logic, which modifies and inspects the device’s system state.
We experiment with baseline agents to test AndroidWorld and provide initial results on the benchmark. Our best agent can complete 30.6% of AndroidWorld’s tasks, leaving ample room for future work. Furthermore, we adapt a popular desktop web agent to work on Android, which we find to be less effective on mobile, suggesting future research is needed to achieve universal, cross-platform agents. Finally, we also conduct a robustness analysis, showing that task variations can significantly affect agent performance, demonstrating that without such testing, agent performance metrics may not fully reflect practical challenges. AndroidWorld and the experiments in this paper are available at https://github.com/google-research/android_world.
@@ 1 Introduction
Autonomous agents that interpret natural language instructions and operate computing devices can provide enormous value to users by automating repetitive tasks, augmenting human intelligence, and accomplishing complex workflows. However, a key research challenge remains the realistic evaluation of these agents in real-world settings. Despite growing enthusiasm for building autonomous agents [45, 11, 81, 27, 25, 20, 17, 1, 2, 63, 65], most existing approaches for evaluation compare an agent’s actions at each step to a previously collected human demonstration [11, 45, 71, 78, 37, 76, 69, 32]. Measuring performance in this way can be misleading because when performing tasks online in real environments agents can take multiple paths to solve tasks, environments may behave non-deterministically, and agents can dynamically learn from mistakes to correct their actions [49, 35, 31, 43]. For this reason, online evaluation of agents in realistic environments able to reward task outcome provides a gold standard for evaluation. While there is an emerging body of work to address this need across different environments [84, 27, 12, 29, 66, 5], there is no comprehensive solution for mobile platforms, such as Android, which are used by billions of users and therefore represent environments in which automation agents may be very productively employed. We introduce AndroidWorld to address this.
At its core, AndroidWorld offers a reliable means of obtaining reward signals for tasks performed by agents in realistic mobile environments. Reward signals are quantitative metrics that indicate functional correctness of a task, i.e. is the stated goal achieved? For example, for the task “Send a text message to Jane confirming I’ll be there,” a positive reward indicates that the relevant message has been sent. Unlike simulated environments [54, 50] or games [41, 52, 60, 62, 53, 58], real-world apps and websites do not inherently offer explicit reward signals. While human [45, 81, 43, 26] or LLM-based [9, 82, 36, 13, 38, 43, 20] judges can be employed to reward the outcome of a task, these approaches scale poorly or are not fully reliable, respectively. Alternatively, environments for autonomous agents which provide automated ground-truth rewards for complex workflows have been developed  [73, 84, 27, 66, 5]. We find two problems with these environments. First, they are constrained to desktop computing environments, overlooking the mobile domain, which is of paramount importance given the ubiquity and diversity of mobile devices in the real world. Secondly, they are limited in their real-world diversity and scale. Crucially, unlike in real-world scenarios where conditions and task inputs vary widely, these environments support only static test specifications, meaning that when task parameters deviate, the reward signal is likely to break.
@@
We seek to develop a comprehensive benchmark that addresses the limitations of the existing approaches above for evaluating automation agents in mobile environments. AndroidWorld does this by spanning 20 Android apps on a total of 116 programmatic tasks to provide ground truth-rewards. Unlike existing test environments (with MiniWoB++ [48] being a notable exception), each task in AndroidWorld is dynamically instantiated using randomly-generated parameters, challenging agents with millions of unique task goals and conditions. While MiniWob++ consists of simple, synthetic websites, AndroidWorld leverages actual Android applications. A main challenge that AndroidWorld must address is how to ensure that reward signals are durable when using real-world applications and varying task parameters dynamically. AndroidWorld’s key insight is to leverage the extensive and consistent state management capabilities of the Android operating system, using the same mechanisms that the apps themselves utilize to store and update data.
In addition to providing a comprehensive benchmark, AndroidWorld is lightweight, requiring only 2 GB of memory and 8 GB of disk space, and is designed with convenience in mind. It connects agents to the Android OS by leveraging the Python library AndroidEnv [58] to connect to the freely available Android Emulator. [^1] The Android Emulator is packaged as part of Android Studio, which can be downloaded from https://developer.android.com/studio In addition to the 116 Android tasks, we extend AndroidWorld with web tasks by integrating the MiniWoB++ [48, 34] benchmark into it.
To demonstrate AndroidWorld’s usefulness as a benchmark, we build and release a multi-modal agent, M3A (Multimodal Autonomous Agent for Android), and establish state-of-the-art results on AndroidWorld. We analyze M3A’s performance using both multimodal and text-only input, and we observe that while multimodal perception can improve performance in some cases, it generally does not outperform the text-only approach. On AndroidWorld, M3A achieves a 30.6% success rate, which surpasses that of a web agent adapted for Android but remains significantly lower than the human success rate of 80.0%. In pursuit of building robust UI control agents, our study includes comprehensive tests under varied real-world conditions, demonstrating significant performance variations primarily driven by changes in intent parameters.
Overall, we make the following contributions: (i) the creation of a new, highly diverse and realistic mobile UI control agent environment; (ii) establishment of benchmark performance with a state-of-the-art multimodal agent, and (iii) a careful analysis demonstrating the need to evaluate agents across variable task parameters and conditions due to the inherent stochasticity in both models and environments.
@@ 2 Related Work
Table 1 compares existing evaluation environments for autonomous UI agents.
@@ 2.1 Interactive evaluation environments
Effective evaluation of autonomous agents requires benchmarks that mimic real-world scenarios, but also interactive environments that provide reward signals upon successful task completion [45, 11, 3, 46, 7]. Many existing benchmarking environments target web browsing. MiniWoB++ [48, 35] consists of small, synthetic HTML pages with parameterizable tasks which allow for unlimited task variability. WebShop [73] provides a simulated e-commerce environment, whereas WebArena [84] and VisualWebArena [27] consist of simulated websites across up to six domains. WorkArena [12] consists of 29 tasks for enterprise software. GAIA [40] is a static dataset that tests an agent’s ability to interact with live web environments. MMInA [80] is a multihop and multimodal benchmark designed to evaluate agents for compositional Internet tasks.
Towards building generalist agents that control operating systems (OSes), OSWorld [66] and WindowsAgentArena [5] provide a test suite of tasks for desktop computer use cases and custom execution-based evaluation scripts across across 9 and 11 apps, respectively. In the mobile domain, existing benchmarks are limited and do not capture the diversity of real-world mobile interactions, containing low-complexity tasks or on a limited number of applications. B-MoCA’s [29] evaluation is based on 6 simple tasks (e.g., ”Call 911”, ”turn on airplane mode”) across 4 apps [^2] Based on what reported in the Experiments Section of the B-MoCA manuscript as of October 1st, 2024., validated using regular expressions. Mobile-Env [75] offers task reproducibility limited to 13 task templates for a single app (WikiHow).
While AndroidWorld shares the mobile platform focus of B-MoCA and Mobile-Env, it is more comparable to OSWorld (and WindowsAgentArena, which builds on top of OSWorld) in terms of task complexity and the diversity of interactions it supports. AndroidWorld enhances OSWorld’s approach by dynamically constructing the start states of an agent’s run and varying the task parameters in unlimited ways, thus allowing for a new type of evaluation under varying real-world conditions.
@@
Other studies leverage human evaluation [45, 81, 4] for tasks where automatic evaluation is not available. Lastly, emerging research [43, 20, 67] explores the potential of multimodal models to generalize agent evaluations to new settings, though this area requires further research to achieve accuracy comparable to manually-coded rewards.
AndroidEnv [58] provides a mechanism to manage communication with the Android emulator, similar to Playwright and Selenium for web environments. While AndroidWorld leverages this functionality, it diverges in its reward system. AndroidEnv’s approach requires modifying application source code and implementing task-specific logging statements, making it well-suited for gaming environments with easily verifiable success criteria. In contrast, AndroidWorld implements a non-invasive reward mechanism, allowing it to create a benchmark suite for apps where source code access is unavailable and to reuse validation components across different apps. This approach enables AndroidWorld to cover a broader range of real-world mobile tasks.
2.2 Static datasets for UI automation
@@
Datasets derived from human interactions provide proxy metrics that correlate with real-world agent performance [33, 6, 11, 45]. On mobile platforms, AitW [45], AndroidControl [32], PixelHelp [33], AndroidArena [67], LlamaTouch [77], UGIF [59], and MoTIF [6] consist of demonstrations across Android apps and mobile websites, with screens often represented via accessibility trees. In contrast, desktop web environments typically utilize the DOM for representing website content, with Mind2Web [11], OmniAct [24] and others, across various desktop websites. Mobile-based datasets frequently involve more complex actions, such as scrolling, which are not as useful in DOM-based desktop interactions where the entire action space is readily accessible. Additionally, API-centric datasets like API-Bank [30], ToolTalk [14], and ToolBench [68] assess agents’ capabilities to manipulate computer systems via APIs.
@@ 2.3 Interactive agents
Prior to today’s foundation models, traditional approaches to developing user interface-operating agents primarily used reinforcement learning and behavioral cloning to simulate interactions like mouse clicks and keyboard typing [35, 33, 51, 18, 23]. More recent work tends to leverage off-the-shelf foundational models [55, 56, 57] with in-context learning (ICL) and fine-tuning applied to mobile [45, 22, 61, 69, 78, 4, 79], desktop web [81, 11, 84, 27, 8, 28], and desktop OS [64, 74, 66]. Recent work explores agents that reflect on system state [49, 72, 39] by leveraging exploration, self-evaluation, and retry-capabilities enabling continual learning and adaptation [31, 71, 43, 64, 16, 42].
@@
AndroidArena  13   221 1 Action match/LLM Android (Apps+Web)
LLamaTouch  57   496 1 Screen match Android (Apps+Web)
@@
Mobile-Env    1    13  11.5 regex Android (Apps)
B-MoCA    4      6    1.9 regex Android (Apps+Web)
@@ 3.2 The observation and action space
AndroidWorld provides an interface for agents to receive observations and execute actions on Android. It uses AndroidEnv [58] and the Android Device Bridge to facilitate interaction between Android and the agent. The observation space consists of a full-resolution screenshot and a UI tree representation developed for accessibility purposes. The action space is similar to that which humans use, consisting of gestures (i.e., tapping, long-press, and swiping), typing, and navigation buttons (i.e., go home and go back). In addition to these naturalistic actions, AndroidWorld exposes a limited set of function calling APIs, such as send_text_message, to help agents accomplish goals. Appendix B provides more details on the observation format and action space.
@@ 3.3 Reproducible and parameterized tasks
Figure 2: Annotators performed the tasks assigned to them, assigned a difficulty level (2(a)) and estimated the number of steps required to complete each task (2(b)), using the action space available to an agent. For each task, they selected relevant category tags from a predefined list (2(c)).
AndroidWorld consists of a suite of 116 tasks, spread across 20 diverse applications (see Appendix C for more details). These tasks simulate practical, everyday activities, including note-taking, scheduling appointments, communicating through messaging, and interacting with system utilities. The suite consists of open-source apps and built-in Android system apps, such as Settings and Contacts. As rated by humans, the tasks vary in difficulty, duration, and categories (Figure 2).
To achieve a high degree of reproducibility in real-world scenarios, AndroidWorld precisely controls the OS and app states in several ways. The Android OS is fixed, consisting of a Pixel 6 emulator running Android 13 with a fixed time on the date October 15th, 2023. All applications in AndroidWorld are fully-functional and consists of both open-source apps and OS-level apps included with Android. For the open-source apps, AndroidWorld maintains a constant environment by installing a fixed version of each app, acquired from F-Droid. [^5] https://f-droid.org/ OS-level apps’ versions are determined by the Android OS, which is also fixed. To maintain a reproducible environment, AndroidWorld utilizes apps that do not require login/authentication and can store their application data on device.
In addition to managing the states of apps and operating systems, AndroidWorld precisely defines and controls the state during task execution. Each task has its own unique setup, reward determination logic, and teardown procedures (see Appendix C.2 and C.3 for more details), ensuring a fully reproducible suite of tasks.
Automatic task parameterization is a critical mechanism, unique to AndroidWorld, to evaluate agents on a much larger and more realistic suite of tasks than current benchmarks support. Achieving this requires significantly more effort than randomly generating new task parameters because it involves developing evaluation logic that remains valid across different task instantiations. It is exactly through its careful state management that in addition to reproducibility AndroidWorld ensures that the reward mechanisms function correctly. Task parameters, initialized randomly at the start of each task based on a controlled random seed, dictate the initial state and influence reward outcomes. Similarly to MiniWoB++ [48, 34], AndroidWorld consists of a practically infinite set of varying initial conditions and success criteria.
This approach provides more granular analyses of agents’ adaptability — a vital attribute for real-world deployment. Beyond testing agent robustness, the dynamic construction of tasks supports the use of online learning methodologies, particularly reinforcement learning [48, 34, 23, 18]. It also simplifies the generation of distinct train/test datasets, facilitating supervised learning experiments [23, 47, 15].